Express the function g: R+-R+ is defined by , g(X) =1/√x+√x, as a composition of three function
1
Expert's answer
2020-02-27T12:13:05-0500

"g(x)=\frac{1}{\sqrt{x + \sqrt{x}}}, \quad g: \mathbb{R}^+ \to \mathbb{R}^+"


Let "f(x)=\sqrt{x}, h(x)=\sqrt{x^2+x}, \text{ and } k(x) = \frac{1}{x}."


Determine "h\circ f."

"h\circ f = h(f(x)) = \sqrt{(f(x))^2 + f(x)} = \sqrt{(\sqrt{x})^2 + \sqrt{x}} \\
\qquad = \sqrt{x + \sqrt{x}}"


Now, determine "k\circ(h\circ f)."


"k\circ(h\circ f) = k(h(f(x))) = \frac{1}{h(f(x))} = \frac{1}{\sqrt{x + \sqrt{x}} }"


Therefore, "g(x) = k\circ(h\circ f), \text{ where } f(x)=\sqrt{x}, h(x)=\sqrt{x^2+x}, 
\text{ and } k(x) = \frac{1}{x}."